To obtain the data needed to plot a titration curve for the titration of a weak acid with a strong base, a chemist used 50.00 mL of 0.10mol/L of citric acid with 0.10 mol/L NaOH(aq). Calculate the pH of the resulting solution at the equivalent point.
1
Expert's answer
2020-06-10T07:37:35-0400

The citric acid is a weak organic tribasic acid. It's chemical formulae can be written for simplicity as AH3. Therefore, the reaction of the titration of the citric acid by sodium hydroxide is:

AH3 + 3NaOH "\rightarrow" Na3A + 3H2O.

According to the stoichiometric coefficients, the number of the moles of NaOH added and the number of the moles of citric acid in the sample relate as:

"\frac{n(NaOH)}{3} = n(AH_3)" .

Therefore, at the equivalence point, the volume of NaOH added is:

"V = \frac{n}{c} = \frac{3c(AH_3)V(AH_3)}{c(NaOH)} = \frac{3·0.10·50.00}{0.10} = 150.00" mL.

The concentration of sodium citrate at the equivalence point is:

"c(Na_3A) = \frac{c(AH_3)V(AH_3)}{V_{tot}} = \frac{0.10·50.00}{150.00+50.00} = 0.025" mol/L.

The citrate anion partly reacts with water:

A3- + H2O "\rightarrow" HA2- + OH-.

According to this reaction, OH- anions are released. The equilibrium constant of this reaction, "K_b" , relates to the third step acidity constant of citric acid as:

"K_b = \frac{K_w}{K_{a3}} = 2.5·10^{-8}" .

As there is two orders difference between "K_{a2}" and "K_{a3}", we can consider only this reaction (the third step of dissociation).

The concentration of OH- anions is therefore:

"K_b = \frac{[HA^{2-}][OH^-]}{[A^{3-}]}" .

The equilibrium concentration of OH- anions equals the equilibrium concentration of HA2- anions. The equilibrium concentration of the citrate anion is (0.025 - x), where x is the equilibrium concentration of OH-:

"(0.025-x)\u00b72.5\u00b710^{-8} = x^2"

"x = 2.5\u00b710^{-5}"

"[OH^-] = 2.5\u00b710^{-5}"

"pOH = 4.6"

"pH = 14-pOH = 9.4"

Answer: the pH of the resulting solution at the equivalence point is 9.4.